Instrument Log 

CSPHOT        C1                  CF Lamont, OK

Date

Technician

Comments

7/1/16

CAM

Turned cloud mode on at 2114 GMT.

6/23/16

JEM

From DQPR #5172:  Aeronet reported the data received has measured 0 signal since it was set up. It might be an instrument problem since it is sunny today. Please check whether PARK, GOSUN, and TRACK functioned correctly. Instrument alignment was off approximately 15 degrees. The instrument would not find the sun with the "Park", "GoSun" and "Track" sequence.  Loosened the mounting bolts and rotated the instrument to the proper position.  Performed "Park", "GoSun" and "Track" again.  This time the instrument tracked the sun as expected.  RTS @2000 GMT.

6/22/16

KT

CSPHOT #620 came back from being repaired and needed to be installed. At 1830 GMT installed CSPHOT #620. At 1930 GMT powered on the instrument and changed the Date and Time. Contacted David Swank and informed him that the instrument was powered on and ready to run and asked if he could turn collections back on. At this time system is online and operational. RTS at 1945 GMT. System Electronics SN# 1003049/ WD# 59463, Sensor Head WD# 79709.

4/1/16

KT

Work order requested that CSPHOT #620 system be removed and sent back for repairs. At 1424 GMT stopped the program and powered down the instrument. There are no spare CSPHOT instruments to use as a replacement until unit #620 is returned to SGP. Robotic Arm SN# 1005-093 WD# 79716, Electronics Enclosure #620, WD# 59463, Photometer #620 WD# 79709. RTS at 1500 GMT.

3/2/2016

MS

A DQPR 5172 has opened up and stated that the CSPHOT cables and collimator tubes needs looked at.  At 1600 GMT put the instrument into park mode to inspect the collimator tubes and found no obstruction.  Also checked the cables as well and found nothing during this inspection.  Put collimator tubes back on and took instrument out of park mode, then put into Auto mode.   RTS 1615 GMT.

1/27/2016

MS

Instrument is down as of 1601 GMT, 1/26/2016 for a scheduled Virtual Computer Host upgrade.  Data will be lost until upgrade is complete.

12/15/15

MS

Computer needed to be restarted following scheduled software and security patch updates.  At 1404 GMT computers were updated and rebooted by the IT department, some data might be lost.

12/11/15

MS

Plot from 12/2/2015 showing data suggestive of a tube obstruction.  At 1640 GMT the instrument was put into Park Mode, then cleaned out the collimator tubes. Did not see any debris in tubes during cleaning. RTS at 1655 GMT.

11/17/2015

MS

A 7 hour power outage starting at 0700 GMT and power was returned at 1435 GMT, 11/17/2015. Instrument was down. During that time some data might be lost?  Started program back up and RTS at 1520 GMT, 11/17/2015.

10/23/15

KT

DQPR-4926 stated that the CSPHOT Collimator tubes were obstructed. At 1734 GMT took instrument out of Auto mode and placed instrument into Park. Removed the Collimator tubes and inspected for obstructions. Could not visually see any obstructions but did clean tubes. At 1740 GMT placed Collimator tubes back on instrument, took instrument out of Park mode and placed into Auto mode. RTS at 1749 GMT.

10/22/2015

MS

Instrument had wrong date and time displayed, also stuck on waiting for port com 3.  Restarted instrument twice with no success between 1324 GMT and 1328 GMT.  Continued with PM rounds, while cleaning the collimator tubes notice that the CSPHOT was displaying the correct time and date at 1435 GMT.

10/16/2015

MS

Responded to a DQPR, spider web in collimator tube, put instrument in park mode at 1530 GMT and removed spider web.  RTS at 1545 GMT.

10/16/2015

MS

Program had wrong date displayed, stopped and restarted program at 1315 GMT  RTS at 1330 GMT.

10/8/15

KT

Cimel_http_connection.exe was not working. Removed the Cimel controller box and brought back to the Electronics Repair Lab and performed testing. Connected smaller data cable to the controller box and to the computer and was able to get the data files and didn't experience the error message. Took the Cimel controller box and placed back into the enclosure, relocated the moxa network hub into the CSPHOT enclosure box and ran a network cable from the enclosure box down to the network hub inside the Optical Trailer. After everything was all connected restarted the program and system appears to be online and operational without any error messages. RTS at 1550 GMT.

10/5/15

KT

SDS reported no file from the instrument. At 1336 GMT the program had stopped working. Cimel_http_connect.exe has stopped working error message was on the screen. Closed the error window and restarted the program. Program came online and displayed waiting for comport 4. RTS at 1351 GMT.

9/28/15

KT

SDS reported no data from the instrument. At 1325 GMT found the CSPHOT electronics had no display. Opened up the electronics enclosure and re-seated connectors. Closed the enclosure and was able to get the display. Reset the time and date. System is online and operational. RTS at 1340 GMT.

9/24/15

KT

We received CSPHOT #620 back from being repaired and tested and the unit needed to be changed out with #622. At 1754 GMT stopped the program and powered down unit #622. Removed and replaced CSPHOT units. At 1840 GMT powered unit #620 on. At 1845 GMT system was online and operational. RTS at 1850 GMT.

*Old - #622 SN# 1003052/UA3902 WD#59464.

*New - #620 SN# 1003049/UA389 no WD#.

9/16/15

KT

The program was not running. At 1248 GMT closed the window and restarted the program. RTS at 1303 GMT.

9/8/15

KT

The computer program was not running. At 1226 GMT started the computer program. At 1233 GMT the computer came online and began normal operations. RTS at 1241 GMT.

7/29/15

JEM

The newly installed CIMEL is having communications troubles that are causing the data to not be collected.  Mentor requests the CIMEL be swapped back with the old one.  R/R *CIMEL control box, head, sensor cable and interface.  Verified parameters matched removed unit and restarted.    RTS at 1530 GMT.

 

*Old 620: Controller S/N - 1003049/UA3892, WD59463  Head S/N - 1003050(T804NE), WD79709

 

*New 622: Controller S/N - 1003052/UA3902, WD59464  Head: S/N - 1003053(T811NE) WD79711

7/24/15

JEM

The CSPHOT is having trouble reading the header information again and is not collecting any data.  Removed the controller again and again there is a very strong odor of overheated electronics.  Tried reseating all of the IC chips again and also checked for any bad solder joints, but did not find any.  There is a PNP Silicon Power Transistor that has obviously been hot before and has been replaced.  This may be the source of the smell and could also be the problem.  After reinstalling the controller, the header errors continued.

7/23/15

JEM

The CSPHOT has not been collecting data because it is having errors when reading the header information.  Also when it tries to transfer data the data/time information says 00:00:2015, 00:00:00, but on the instrument itself the time is correct.  This is probably due to the header problems.  Tried using a different serial port to communicate with the CSPHOT.  The same problems occurred.  Changed the junction board on the front of the controller box.  Still had the same problems.  Checked the cable for any problems or corrosion and no problems were found.  Removed the controller and reseated all IC chips on the boards.  When the board was removed there was a strong odor of overheated electronics.  Returned the controller to the system and powered it up.  Restarted the collection software and it started to properly transfer data for approx. the last 24 hours.  RTS @2000 GMT.

7/17/15

PAD

DQPR 4767 indicates a curve in pwv data and an InGaAs/Silicon discrepancy error.  Unit 620 has arrived on site and Mentor request unit swapped out with unit 622.  R/R *CIMEL control box, head, sensor cable and interface.  Verified parameters matched removed unit and restarted. 

RTS at 1840 GMT.

 

*Old - 622: Controller CE318NEBS9_C, SN 1003052/UA3902, WD59464.  Head: CE318NEBS9_T, SN 1003053(T811NE),  No WD 

 

*New - 620: Controller CE318NEBS9_C, SN 1003049/UA3892, WD59463.  Head CE318NEBS9_T, SN 1003050(T804NE), WD79709. 

6/19/15

JEM

DQPR #4767 requests that the CSPHOT tubes be checked for an obstruction.  Checked the tubes @1827 GMT and found a few small webs down in the center sections of the tubes.  Removed the webs and returned to service @1845 GMT.

6/8/2015

MS

Put CSPHOT into park at 1932 GMT to inspect collimater tubes.  DQPR 4767 has opened for this situation.  Found a single spider web going across the window on one of the collimator tubes.  Cleaned in and outside of both collimator tubes at 1932 GMT.  RTS 1947 GMT.

4/29/15

MS

CSPHOT stopped collecting data.  Closed out program and restarted CSPHOT at 2001 GMT.  RTS status O.K.

1/27/15

JEM

According to DQPR #4519 , the collimator tube appears to be obstructed.  Took the CSPHOT out of auto mode and inspected the tubes for any obstruction.  No obstruction was found.  Returned the CSPHOT to auto mode and RTS @2145 GMT.

11/26/14

PD

Mentor requests *Cimel units swapped.  Powered down and removed unit 402 sensor, controller, sensor cable and collimator tubes.  Installed unit 622.  Old Controller Type CE318NVBS5_C, SN 0605115/UA3436, WD 79719.   Old Sensor Type CE318NVBS5_T, SN 0605116 (T562N), WD48464.  New Controller type CE318NEBS9_C, SN 1003052/UA3902, WD59646.  New Sensor CE318NEBS9_T, SN 1003053 (T811NE).  Verified instrument aligned with sun after install.  RTS at 2020 GMT.

*Old - 402  *New - 622

10/10/14

JEM

In DQPR #4304 the mentor indicates that the 1640 channel is no longer following the solar curve, but it still off and requests the tubes be inspected again.  Checked the tubes again at 2030 GMT and found what looked like a small egg sack and webs on the wall of the tube very near the sensor end. It was very hard to see unless the tube and light were at the right angle. Removed the tube and took it to the lab to clean. With the help of a bright light and some dental picks, I believe I was able to remove all of the webs and egg sack. Returned the tubes to the CSPHOT and RTS @2050 GMT.

10/07/14

JEM

DQPR #4304 indicates a possible blockage in the collimator tubes.  Checked the collimator tubes @2030 GMT and found a good amount of spider web inside.  Most of the web was around the walls of the tubes and did not create a complete blockage.  There was some spider web in across the outside of the tube opening along with a spider.  Removed all webs and the spider @2040 GMT.  RTS @2045 GMT.

7/16/14

CAM

No data had been collected from CSPHOT for 43 hours.  Remote status check found cimel_http_connect was running but did not appear to be transferring files.  Closed and restarted program at 2000 GMT.  Today’s data file was found and uploaded.  RTS status O.K.

6/9/14

CAM

No data had been collected from CSPHOT since 0006 GMT on 6/8.  Remote status check found cimel_http_connect was running but did not appear to be transferring files.  Closed and restarted program at 1221 GMT.  Today’s data file was found and uploaded.  RTS status O.K.

4/21/14

CAM

No data had been collected from CSPHOT for 36 hours.  Remote status check found cimel_http_connect was running but did not appear to be transferring files.  Closed and restarted program at 1212 GMT.  Today’s data file was found and uploaded.  RTS status O.K.

4/3/14

PAD

Cimel clock is 4 sec slow.  Adjusted clock at 1449 GMT.  RTS at 1504 GMT.

10/16/13

CAM

No files had been collected from the CSPHOT since 10/14.  Remote status check found cimmel_http_connect to COM3 was not running.  Started program and data files were uploaded.  RTS at 1225 GMT.

10/3/13

JEM

The CSPHOT has not collected any data since the new instrument was installed on 9/23/13.  Troubleshooting determined a faulty MOXA device.  Removed the MOXA 5450I and replaced it with a MOXA W2250.  Configured the computer to map COM3 to Port1 of the MOXA.  Restarted the CIMEL Connect software and the CSPHOT began transferring data.  RTS @1600 GMT.

9/23/13

KT

Work order requests that the CSPHOT be removed and replaced. At 1640 GMT powered down the *CSPHOT and removed and replaced system. RTS at 1658 GMT. Old CSPHOT Controller #629, MD# CE318NEBS9_C, SN# 1005097/UA3949. Old Sensor Head #629, MD# CE318NEBS9_T, SN# 1005098(T834NE). New CSPHOT Controller #402, MD# CE318NVBS5-C, SN# 0605115/UA3436, Sensor Head #402, MD# CE318NVBS5-T, SN# 0605116(T562N), WD# 48464.

8/5/13

CAM

Last SDS collection was on 8/1/13 at 0006 GMT.  Accessed computer and Com1 box was shown instead of Com 3.  Closed program and tried to start "Cimel_http_connect to Com3" program.  Program would open briefly but then close.  Restarted computer and tried again with same results.  Rebooted computer once more and ran “Cimel_http_connect” and then was able to successfully start “Cimel_http_connect to Com3”.  RTS at 1223 GMT.

7/22/13

KT

SDS reported no data collected from instrument. The computer program was not running and needed to be started. At 1225 GMT started the program. RTS at 1240 GMT. Status OK.

7/2/13

KT

The instrument was not scanning. At 1410 GMT found that the instrument was not in AUTO mode. At 1413 GMT place the instrument in AUTO mode. RTS at 1425 GMT.

7/1/13

JEM

Last SDS collection was on 6/28/13 at 0006 GMT.  The mentor said if the CSPHOT could not be fixed to remove and send in for repair and calibration.  Disconnected all power sources for approximately 15 minutes.  Reconnected all power sources @2000 GMT.  The screen came on but the date/time was 6/28/13 01:30.  Set the time to current time @2002 GMT.  Verified other settings were correct.  Using the display buttons ran "Go Sun", "Track" and "Park" several times with no trouble.  Talked to mentor and decided to leave the CSPHOT in place until we receive a replacement or until it fails again.  RTS @2010 GMT.

7/1/13

CAM

Last SDS collection was on 6/28/13 at 0006 GMT.  Logged into computer remotely and found there were no current data files.  Checked instrument and found power was being supplied to it, but the display was blank and would not respond when buttons are pressed.  Suspect a memory chip failure as symptoms are the same as when the chip failed recently.  At that time the chip was swapped with another one on the board at the vendor’s request.  Notified mentor of the situation and waiting for response on how to obtain replacement chip.  Instrument remains down at this time.

5/8/13

CAM/JEM

 

Instrument was removed on 4/17 so the wooden decking and stairs could be replaced with metal.  Construction work is complete and instrument needs to be reinstalled.  Reinstalled CSPHOT.  A new data cable was built in the lab to replace the old cable.  Leveled instrument and set date/time at 1920 GMT.  Confirmed alignment with sun and RTS at 1925 GMT.

4/17/13

CAM/JEM

The optical trailer wooden decking and stairs are to be removed and replaced with metal decking and stairs.  Need to remove all instruments so removal of old decking can start tomorrow. Stopped CSPHOT program and powered down at 1402 GMT.  Removed instrument, control box and mounting stand as one complete unit and stored inside trailer.  Instrument will remain offline until new deck and stairs are complete. 

3/25/13

CAM

Last SDS collection was on 3/22/13 at 0006 GMT.  Remotely accessed computer and found COM1 box was shown instead of COM3.  Closed program and attempted to start "Cimel_http_connect to COM3" program.  Program window would open briefly and then close. A new folder was also created on the desktop with each attempt to open the program.  Folders were named dir_com#, where # incremented by one with each attempt.  Rebooted virtual PC with no change in status.  After several attempts at opening COM3 program, it finally stayed open and showed data files ready to transfer.  Checked data directory and confirmed a new file was started. Deleted folders that had been created on the desktop.  RTS at 1223 GMT.

3/14/13

PAD

CIMEL clock is 4 seconds fast.  Reset clock at 1620 GMT.

2/28/13

PAD

CIMEL clock is 3 seconds fast.  Reset clock at 1529 GMT.

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Powered optical cluster UPS off at 1730 GMT.  Power was restored at 1920 GMT.  Restored UPS power at 1945 GMT. Instrument RTS status O.K and data collections resumed once SDS operations were restored.

2/25/13

PAD

Last SDS collection was on 2/22/13 at 0200 GMT.  Accessed computer and Com1 box was shown instead of Com 3.  Closed program and started "Cimel_http_connect to Com3" program.  Computer connected with instrument.  RTS at 1318 GMT. 

2/19/13

PAD

CIMEL clock is 5 seconds slow.  Adjusted clock at 1605 GMT.

2/5/13

PAD

CIMEL clock is 4 seconds fast.  Reset clock at 1545 GMT.

1/7/13

CAM

No files had been collected from the CSPHOT since 1/3.  Remote status check found cimmel_http_connect to COM3 was not running.  Started program and data files were uploaded.  RTS at 1400 GMT.

12/17/12

CAM

No files had been collected from the CSPHOT since 12/13.  Remote status check found cimmel_http_connect to COM3 was not running.  Started program and data files were uploaded.  RTS at 1310 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  CSPHOT computer was gracefully shut down at 1352 GMT in anticipation of power outage.  Power was restored at 2115 GMT after electrical upgrades were complete.  Restarted CSPHOT computer and RTS at 2125 GMT.  Note CSPHOT should have continued to run on battery during power outage and all data should be recovered.

11/14/12

PAD/KT

DQPR 3582 states no data has been collected since 11/7/12.  CIMEL #99 was replaced with unit #628 on 11/7/12.  At 1655 GMT, checked settings on CIMEL and found Link to PC was set to NO.  Changed setting to YES.  Data is now being sent to virtual computer.  RTS at 1710 GMT.

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  CSPHOT auto-started status O.K.

11/7/12

KT

Instrument mentor requested that the CSPHOT be changed out and the old one sent back for calibrations. At 1900 GMT stopped the computer. Removed and replaced *CSPHOT instrument and electronics. At 2000 GMT powered on, set time and date. RTS at 2007 GMT.

*Old #98 SN# 9905-938 176(UA1721),WD# 35516.

*New #629 SN# 1005097/UA3949

9/18/12

PAD

CIMEL clock is 4 seconds slow.  Adjusted clock at 1349 GMT.

9/12/12

PAD

CIMEL clock is 4 seconds slow.  Reset clock at 1508 GMT.

8/30/12

PAD

CIMEL clock is 3 seconds slow.  Adjusted clock at 1409 GMT.

8/23/12

PAD

CIMEL clock is 4 seconds slow.  Adjusted clock at 1517 GMT.

8/17/12

PAD

CIMEL was left out of Auto mode during PM rounds on 8/16/12.  Placed instrument in Auto mode during PM rounds.  RTS at 1504 GMT.

6/18/12

PAD

Last SDS collection was on 6/17/12 at 2300 GMT.  At 1307 GMT, remotely accessed computer and found CIMEL program was not running.  Started program and computer connected with instrument.  All data should have been stored on instrument and transferred to computer with no loss of data.  RTS at 1322 GMT.

6/13/12

PAD

CIMEL clock is 6 seconds slow.  Adjusted clock at 1454 GMT.

5/31/12

PAD

CIMEL clock is 7 seconds slow.  Adjusted clock at 1551 GMT.

5/15/12

PAD

CIMEL clock is 4 seconds slow.  Adjusted clock at 1458 GMT.

5/7/12

PAD

CIMEL clock is 5 seconds slow.  Adjusted clock at 1540 GMT.

4/25/12

PAD

CIMEL clock is 4 seconds slow.  Reset clock at 1411 GMT.

4/17/12

PAD

CIMEL clock is 5 seconds slow.  Adjusted clock at 1455 GMT.

4/9/12

PAD

CIMEL clock is 4 seconds slow.  Adjusted clock at 1532 GMT. 

4/2/12

PAD

CIMEL clock is 3 seconds slow.  Adjusted clock at 1427 GMT.

4/2/12

PAD

Mentor requests cloud mode turned on CIMEL.  Changed BCL-SKY setting from No to Yes at 1425 GMT.  RTS at 1440 GMT.

3/27/12

PAD

CIMEL clock is 4 seconds slow.  Adjusted at 1755 GMT.

3/19/12

SPD

CSPHOT program was not running.  Restarted program @ 1345 GMT. Program appears to be running. RTS @ 1350 GMT.

3/12/12

PAD

CIMEL clock is 5 seconds slow.  Adjusted clock at 1456 GMT.

3/5/12

PAD

CIMEL clock is 4 seconds slow.  Reset clock at 1619 GMT.

2/29/12

PAD

CIMEL clock is 4 seconds slow.  Reset clock at 1632 GMT.

2/23/12

PAD

CIMEL clock is 4 sec slow.  Reset clock at 1715 GMT.

2/14/12

PAD

CIMEL clock is 7 seconds slow.  Adjusted clock at 1538 GMT.

2/6/12

PAD

CIMEL clock is 5 seconds slow.  Adjusted clock at 1513 GMT.

1/30/12

PAD

CIMEL clock is 6 seconds slow.  Adjusted clock at 1608 GMT.

1/23/12

PAD

CIMEL clock is 6 seconds slow.  Adjusted clock at 1529 GMT.

1/17/12

PAD

CIMEL clock is 6 seconds slow.  Reset clock at 1630 GMT.

1/10/12

PAD

CIMEL clock is 4 seconds slow.  Adjusted clock at 1554 GMT.

1/3/12

PAD

CIMEL clock is 5 seconds slow.  Adjusted clock at 1633 GMT.

12/28/11

PAD

Mentor requests cloud mode turned off in CIMEL parameters.  Changed BCL-SKY from Yes to No at 1607 GMT.  RTS at 1622 GMT.

12/27/11

PAD

CIMEL clock is 5 seconds slow.  Adjusted clock at 1422 GMT.

12/21/11

PAD

CF received rain on 12/19 and 12/20/11 and has not produced data since 12/19 at 2026 GMT.  CIMEL shows a reading of "1" for rain sensor indicating rain.  Plugged in spare rain sensor and did not change rain sensor reading.  R/R CIMEL interface located on front of CIMEL box and rain indicator changed to "0".  Placed instrument in auto mode and instrument started taking a measurement shortly after.  RTS at 2041 GMT.

12/15/11

PAD

CIMEL clock is 2 seconds slow.  Adjusted clock at 1608 GMT.

12/12/11

PAD

CIMEL clock is 4 seconds slow.  Reset clock at 1550 GMT.

12/7/11

PAD

CIMEL clock is six seconds slow.  Adjusted clock at 1609 GMT.

12/1/11

PAD

CIMEL clock is 6 sec slow.  Adjusted clock at 1937 GMT.

11/21/11

PAD

CIMEL clock is 5 seconds slow.  Adjusted clock at 1559 GMT.

11/14/11

PAD

CIMEL clock is 4 seconds slow.  Adjusted clock at 1607 GMT.

10/31/11

PAD

CIMEL clock is 4 seconds slow.  Reset clock at 1606 GMT.

10/11/11

PAD

Last SDS collection was on 10/10/11 at 2300 GMT.  Computer shows last communication or data received was on 10/8/11 at 2151 GMT.  Cimel_http_connect program is running.  At 1441 GMT, verified instrument still had power.  Instrument showed no problems when put in manual mode and daily PM was performed.  At 1449 GMT, rebooted virtual computer and restarted Cimel program.  Later verified instrument is sending current data to virtual computer.  RTS at 1456 GMT.

10/11/11

PAD

CIMEL clock is 5 sec slow.  Adjusted clock at 1442 GMT.

10/3/11

PAD

CIMEL clock is 5 sec slow.  Adjusted clock at 1408 GMT.

9/22/11

PAD

Adjusted Cimel clock +4 sec at 1516 GMT.

9/14/11

PAD

Adjusted CIMEL clock +3 sec at 1513 GMT.

9/8/11

PAD

CIMEL clock is 4 sec slow.  Adjusted clock at 1713 GMT.

8/29/11

PAD

Adjusted CIMEL clock +4 seconds at 1526 GMT.  RTS at 1541 GMT.

8/9/11

PAD

CF lost power on 8/8/11 at 2247 GMT.  Power was restored on 8/9/11 at 0211 GMT.  At 1259 GMT, found cimel_http_connect program was not running.  Started program and virtual computer connected with instrument.  RTS at 1314 GMT.

7/13/11

PAD

CSPHOT has not communicated with SDS computer since 7/3/11 after storms move through CF area.  Extensive TS has failed to correct the issue.  Unit 628 is due to be replaced with calibrated instrument and this will hopefully verify or eliminate instrument being cause of comm failure.  At 2003 GMT, R/R *CSPHOT sensor and control box.  After entering correct lat/long coordinates, instrument will align and track sun.  Instrument will now communicate with virtual computer.  *Old Unit 628, Type CE318NEB59_C, No 1005091/UA3947, WD59530.  Head Type CE318NEBS9-T, No 100592(T828NE).  *New Type CE 318-1C, No 9905-938, WD35516, Head WD35515.

6/22/11

PAD

Cimel clock is 3 sec slow.  Adjusted clock at 1445 GMT.

4/21/11

PAD

Remotely accessed virtual computer and program was not running.  Started Cimel_http. connect program.  Computer is now communicating with instrument.  RTS at 1253 GMT.

4/12/11

PAD

Changed BCL_SKY, cloud mode, to yes in parameter settings as per mentor’s request at 2104 GMT.

3/15/11

PAD

CIMEL clock is 4 seconds fast.  Reset clock at 1436 GMT.

1/26/11

PAD

CIMEL clock is 4 seconds fast.  Reset clock at 1624 GMT.

1/18/11

PAD

CIMEL clock is 6 seconds fast.  Adjusted at 1647 GMT.

12/28/10

PAD

CIMEL clock is 5 seconds fast.  Reset clock at 1558 GMT.

12/15/10

PAD

CIMEL clock is 6 seconds fast.  Reset clock at 1624 GMT.

12/1/10

CAM

Site power needs to be shut off to install the transformer and run power to the new office building.  Instrument shut down at 1530 GMT.  Restarted after power was restored and RTS at 1836 GMT.

11/30/10

PAD

CIMEL clock is 8 seconds fast.  Reset clock at 1604 GMT.

10/19/10

MK

As indicated in DQPR 2519, some days the AOT data plots are showing a curve. The mentor has requested unit #99 be replaced with unit # 628. Powered off at 1415 GMT, removed #99 collimator, sensor, robot and electronic control box. Installed #628 new electronic control box, robot, sensor and collimator . Leveled robot assembly, set time, date and parameters. Old #99 electronic box 0305924/UA3080, head 9707-174, robot 9905-946. New #628 electronic box 1005-091 / UA3947, Head 1005-092 / T 828 NE, Robot 1005-093.

9/20/10

PAD

Instrument robot is parking 20-30 degrees off home position in elevation.   Mentor suspects failed limit switch in robot arm.  Removed robot arm and installed replacement.  New holes had to be drilled and tapped for leveling legs on base.  Leveled robot and connected sensor.  Sensor will now park in correct position and align with sun in "Track" mode.  RTS at 2037 GMT.  Old Model CE 318-1 M.  SN 9707-178.  New Model CE 318.  SN 9905-946.

8/24/10

PAD

CIMEL clock is 7 seconds slow.  Adjusted at 1348 GMT.

8/13/10

PAD

During PM rounds during alignment check, instrument went to approximate alignment position after selecting "go sun" mode but would not align with sun after selecting "track" mode.  Parked instrument and tried again with same results.  Rechecked sensor alignment window for obstruction with none found.  Disconnected/ reconnected sensor cable and repeated alignment check with same results.  On fourth attempt, sensor moved into correct alignment with sun.  RTS at 1406 GMT.

7/30/10

PAD

Mentor request cloud mode setting checked on instrument.  At 1428 GMT, changed BCL_SKY from “No” to “Yes” as per mentor’s instruction.  RTS at 1443 GMT.

7/26/10

PAD

Mentor requests unit #402 removed and #99 installed.  Removed #402 controller, cable, sensor and collimator tubes.  Installed unit #99 and reconnected all components.  Set date, time, and checked alignment with sun.  Verified instrument is communicating with virtual computer.  RTS at 1849 GMT.  *Old Unit #402, (Type CE318NVBS5-C.  No 0605115/UA3436.  WD 48464).  *New unit #99, (Type CE318-1.  No 0305924/UA3080.  WD 24960, may be 84960-tag is weathered).

4/7/10

TMG

BCR-01623 requests that new data collection and transfer software be installed on the Cimel computer. 

Removed ASTP Startup link from SDS startup folder.

Disabled ASTP Scheduled tasks.

Installed cimel_http_connect.exe into ftphome directory.

Placed cimel_http_connect.exe com3 shortcut into SDS startup folder. Note: Com3 at the end of the command line is required for the connection to the rocket port on com3. Default is com1.

Note: Data will be written to the following Directory. C:\ftphome\dir_com3

There is not a graphic display now. The only window on the screen will be a scrolling dos window with values displayed.

RTS at 1908 GMT.

10/13/09

PAD

CIMEL clock is 12 seconds fast.  Adjusted at 1502 GMT.

8/31/09

PAD

CIMEL clock is 10 seconds fast.  Reset clock at 1436 GMT.

8/26/09

PAD

CIMEL clock is 7 seconds fast.  Reset clock at 1536 GMT. 

8/24/09

PAD

DQPR 2192 states "low external battery" issue with instrument.  New battery was installed on 8/11/109 and had 12.7 V output at that time.  Output voltage is now down to 11.6V DC.  Power supply for PS-12 controller is 12V DC.  It measures 14.2V DC input voltage at PS-12.  R/R power supply with 20V AC at 1851 GMT.  Input voltage now measures 16.4V AC at PS-12 and output voltage is 12.07V DC and slowly increasing.  Will continue to monitor.  RTS at 1911 GMT. 

8/11/09

PAD

DQPR 2192 states "low external battery" and indicates approximately 11V input to CIMEL controller.  Battery tested 11.1V and input to PS-12 controller tested 14.2V.  Disconnected battery from CIMEL and voltage was unchanged.  R/R battery on PS-12.  New battery shows 12.7V.  Will monitor voltage over next couple of days.  RTS at 2041 GMT.

7/15/09

PAD

DQPR 2192 states that sometimes there are zero signals at almucantar measurements.  12V battery was installed on 6/23/09 for testing purposes and removed on 7/13/09.  Problem only happened late in the day while on battery power.  Aeronet indicates the sensor head cable is most likely restricting robotic arm from full range of movement during that time.  Problem returned in the morning when battery was removed and instrument was reconnected with power supply.  Aeronet suggested installing a battery with regulator.  At 2001 GMT, shut down instrument and installed Campbell Scientific, model PS 12 / SN 5216.  Voltage is now 13.9 input and 11.2 V output to controller.  Adjusted sensor head cable to allow full range of robotic arm.  RTS at 2031 GMT.

7/13/09

PAD

Observed collimator tubes looking straight up at 1410 GMT.  Battery that had been placed on instrument for testing shows 8.4V.  Removed battery and connected power supply for robotic arm.  Power supply shows 14.7V.  Instrument will now align with sun.  RTS at 1425 GMT.

6/23/09

MK

There may be a problem with the 12v power supply. Aeronet has suggested replacing the power supply with a 12 v battery for a few days.  Checked the power supply, it is currently measuring 14.7. Disconnected power supply and connected a battery currently measuring 12.7 volts. Data will need to be observed for any changes in the almuncantor data.

5/4/09

PAD

CIMEL clock is 6 seconds fast.  Reset clock at 1410 GMT.

4/1/09

PAD

Instrument is being tested with virtual computer.  At 1425 GMT, observed CIMEL in park position but status shows:  Scenario running.  Unable to put in manual mode due to scenario running even after a couple of minutes.  Disconnected/ reconnected sensor head and cleaned windows.  Manually put sensor in park position and heard CIMEL activate.  Sensor moved into correct alignment with the sun.  Will continue to monitor.

3/31/09

KR

CSPHOT moved to virtual collector machine for testing purposes.  Instrument will still be operating normally, but data will be collected through the virtual machine.

2/5/09

TMG

Scheduled network backup of CSPHOT computer failed.  Also unable to connect via RAdmin.  Shut down Cimel collection software and rebooted computer at 2058 GMT.  Restarted and RTS status O.K. at 2100 GMT

12/31/08

PAD

CIMEL clock is 8 seconds fast.  Reset clock at 1528 GMT.

12/9/08

PAD

Computer has power but is not active.  CF was without electricity between 1014     and 1222 GMT.  Restarted computer and program restarted correctly and is communicating with CIMEL.  RTS at 1609 GMT.

10/1/08

CAM

Data collections are failing and computer cannot be accessed remotely.  Found that RAdmin was not running on the computer.  Tried to start RAdmin but it could not be started.  Rebooted computer at 1744 GMT.  RAdmin started on bootup and computer is now accessible remotely.  SDS personnel will look into collection issue.  RTS at 1750 GMT.

7/24/08

PAD

DQPR 1996 indicates "AOD and inversion data products are questionable.  There is an unexpected distinct cosine curve in the AOD data.  Controller is also drifting in time.  Time drift is happening more frequently and can be off by as much as 30 min overnight.  Removed *CIMEL Electronique sensor and **control box.  Instrument has been taken to shipping dept to be sent for repairs and calibration.  RTS at 1519 GMT.  *Old - 9707-174  (unit 99, WD35512, model CE 318-1-T)  **Old - 0305924/ UA 3080  (WD24960, model CE 3-18-1) 

7/18/08

PAD

CIMEL clock is 6 min. 9 sec. slow.  Reset clock at 1458 GMT.

7/16/08

PAD

CIMEL clock is 3 min. 17 sec. slow.  Reset clock at 1455 GMT. 

7/2/08

PAD

CIMEL clock is 20 sec. slow.  Reset clock at 1450 GMT.  RTS at 1505 GMT.

6/24/08

PAD

CIMEL will not align with sun in track mode and instrument is slightly off level.  Leveled instrument.  Disconnected/ reconnected cable from controller to sensor.  CIMEL will align with sun.  RTS at 1521 GMT.

6/18/08

PAD

CIMEL clock is 1 min. 39 sec. slow.  Adjusted clock at 1606 GMT.

6/17/08

PAD

Cimel clock is 9 min. 32 sec. slow.  Adjusted clock at 1748 GMT.

6/16/08

PAD

Power outages at 0429 GMT and again at 1834 GMT.  Computer was off after both power outages.  Restarted computer at 1433 GMT and again at 2010 GMT.  RTS at 2025 GMT.

6/9/08

PAD

CF experienced extended power outage early on 6/9/08.  Computer is down at this time.  Started computer at 1530 GMT.  RTS at 1545 GMT.

6/4/08

PAD

Cimel clock is 1 min. 7 sec. slow.  Reset clock at 1420 GMT.  RTS at 1435 GMT.

5/20/08

PAD

CIMEL clock is 11 sec. slow.  Reset clock at 1439 GMT.  RTS at 1454 GMT. 

5/16/08

PAD

CIMEL clock is 3 min. 44 sec. slow.  Reset clock at 1456 GMT.  RTS at 1511 GMT.

5/9/08

PAD

CIMEL clock is 7 min. 26 sec. slow.  Reset clock at 1520 GMT.  RTS at 1535 GMT.

4/28/08

PAD

On 4/26/08, CF experienced extended power loss approximately 0000 GMT.   At 1338 GMT, found computer has power but is not running.  Started computer and program started correctly.  RTS at 1353 GMT.

4/14/08

PAD

Sensor did not track sun during PM check.  Disconnected/ reconnected sensor cable.  Sensor will now track sun.  RTS at 1529 GMT.

4/8/08

PAD

Power outage at CF between 0709 and 0910 GMT.  Computer has power but nothing is shown on monitor.  Contacted computer department.  Tim Grove was able to remotely access CSPHOT computer and had to reset monitor settings.  RTS at 1456 GMT.

2/4/08

PAD

Sensor will not align with sun.  Disconnected/ reconnected sensor cable.  Sensor will now align with sun.  RTS at 1625 GMT.

1/14/08

PAD

Sensor failed to align with sun during PM check.  Disconnected/ reconnected sensor cable at head.  Sensor will now track sun.  RTS at 1622 GMT.

12/31/07

PAD

Instrument did not track sun during PM check of alignment.  Disconnected/reconnected sensor cable.  Sensor will now align with sun.  RTS at 1556 GMT.

12/24/07

MK

Due to a power outage the UPS has completely drained the internal batteries and is unable to restart. Reset UPS at 1430 GMT, all instruments in the Optical facility have been restarted and are currently running.

12/19/07

PAD

Display stopped graphing approximately 0030 GMT.  Stopped program and power cycled instrument.  Restarted AERI program.  RTS at 1639 GMT.

12/17/107

PAD

Instrument will not align with sun.  Disconnected/ reconnected sensor head cable and tried alignment.  Instrument will now align with sun.  RTS at 1919 GMT.

7/9/07

PAD

At 1816 GMT, CIMEL clock shows 1814:16.  Reset CIMEL clock at 1819 GMT.

6/22/07

PAD

At 1433 GMT, CIMEL clock shows 1308:13 GMT.  Adjusted CIMEL clock at 1435 GMT.  RTS at 1450 GMT.

6/19/07

PAD

At 1529 GMT, CIMEL clock shows 0533 GMT.  Reset clock at 1533 GMT.  Checked instrument alignment with sun.  RTS at 1543 GMT.

5/18/07

PAD

CIMEL clock is 3 min. 4 sec. slow.  Reset CIMEL clock at 1810 GMT.  RTS at 1825 GMT.

5/14/07

PAD

CIMEL clock is 4 min. 16 sec. slow.  Adjusted  CIMEL clock at 1453 GMT.  RTS at 1508 GMT.

5/3/07

PAD

When CSPHOT robot arm was parked during PM rounds, CIMEL display screen went blank and robot arm did not park.  Pushed green button on display, no results.  Unplugged/ plugged 12V input to CIMEL interface, no results.  Tested input voltage at 14.6V.   Unplugged sensor head from CIMEL and plugged in, no results.  After approximately 5 minutes, screen came back, set robot to park and when I pressed go, screen shows Hello, photo...,  then goes blank.  Repeated above process with no results.  Screen came back in 5 minutes and this time CIMEL accepted park command.  Screen was visible but very faint while robot was parking.  Sensor head is not aligned with sun.  Checked instrument again at 1722 GMT after locating instrument parameters. Robot arm will park and sensor will align with sun.  All parameters verified OK.  RTS at 1737 GMT.

5/2/07

PAD

CIMEL clock is 1 min. 42 sec. slow.  Adjusted clock at 1510 GMT.  RTS at 1525 GMT.

4/30/07

CAM

Christine Chiu notified Site Ops that BCL_SKY value was inadvertently left set to NO when it left GSFC after being repaired.  This value needs to be changed to YES.  Changed BCL_SKY to YES at 1740 GMT.  RTS status O.K.

4/24/07

PAD

Instrument is facing in direction of sun but it is overcast.  Last data collected by computer was at 1449 GMT.  CIMEL battery voltage is 4.7V.  CIMEL has 13.6V input at controller connector.  Plugged in solar panel to controller.  Battery voltage immediately climbed to 4.9V and instrument parked.  Unable to check alignment due to sky conditions at this time.  RTS at 1559 GMT.   

4/23/07

PAD

CIMEL clock is 5 min. 50 sec. slow.  Reset clock at 1441 GMT.

4/17/07

MK

CSPHOT S/N 99 has returned from repair and needs to be installed. Powered off S/N 92 at 1511 GMT, R/R *CSPHOT at 1515, powered back on at 1520 GMT, updated collection file and returned to service at 1530 GMT.   *Old – 92   *New - 99

4/12/07

PAD

Observed instrument facing east but not aligned with sun.  CIMEL shows 4.7V and low battery.  CIMEL interface has 14.6V input.    Removed CIMEL internal battery and charged to 5.1V.  RTS at 1514 GMT.

4/4/07

PAD

Observed instrument facing East but not aligned with sun.  CIMEL internal battery voltage shows 4.7V.  Input voltage to CIMEL shows 14.6 V.  Called instrument mentor.  Advised to remove internal battery and charge to 5.0V.  At 1805, removed CIMEL battery and charged.  Installed battery and checked instrument alignment.  Instrument will track sun.  RTS at 1925 GMT.

3/2/07

mk

The loaner CSPHOT needs to be installed while S/N 99 is out for repairs and upgrades.  Installed *CSPHOT at 1855 GMT.   *New S/N 92.

2/5/07

PAD

Wayne Newcomb request sensor head, controller, and controller/ sensor cable and connector removed and sent for troubleshooting and/ or repairs.  At 2100 GMT, removed *CSPHOT sensor, controller (WD 24960), and cable/ connector.  Took all components plus spare cable/ connector to shipping.  RTS at 2120 GMT.   *Old - 9707-174 (model CE 318-I-T)  (#99)

2/2/07

PAD

Wayne Newcomb request sensor cable and connector panel changed out on instrument.  At 1551 GMT, R/R controller to sensor cable and connector panel.  After installation, verified instrument will track sun.  Notified instrument mentor.  RTS at 1621 GMT.

12/20/06

PAD

CIMEL is sending the same files repeatedly to the computer.  Mentor request change in program setup to solve the problem.  In the ASTPWin program, went to general setup under tools and selected Optimize Transmissions.  RTS at 1559 GMT.

12/5/06

PAD

At 1637 GMT, CIMEL clock shows 1634:49.  Reset clock at 1640 GMT.

9/28/06

PAD

Mentor requests *CIMEL and **sensor removed for calibration.  At 1450 GMT, Mark Klassen and I removed *CIMEL (WD24960, #99), **sensor (WD35512, #99) and CIMEL to sensor cable to send for calibration.

9/25/06

PAD

CIMEL clock is 4 seconds slow.  Reset clock at 1636 GMT.  RTS at 1651 GMT.

9/11/06

PAD

CIMEL clock is 8 seconds slow.  Reset clock at 1525 GMT.  RTS at 1540 GMT.

9/5/06

PAD

CIMEL clock is 6 seconds slow.  Reset clock at 1508 GMT.  RTS at 1523 GMT.

8/30/06

PAD

CIMEL clock is 5 seconds slow.  Reset clock at 1514 GMT.  RTS at 1529 GMT.

8/21/06

PAD

CIMEL clock is 6 min. 30 sec. slow.  Reset clock at 1555 GMT.  RTS at 1610 GMT.

8/17/06

PAD

CIMEL clock is 6 seconds slow.  Reset clock at 1513 GMT. RTS at 1528 GMT.               

8/11/06

PAD

CIMEL clock is 5 seconds slow.  Reset clock at 1511 GMT.  RTS at 1526 GMT.

8/7/06

PAD

CIMEL clock is 5 seconds slow.  Reset clock at 1529 GMT.  RTS at 1544 GMT.

7/31/06

PAD

CIMEL clock is 5 seconds slow.  Reset clock at 1538 GMT.  RTS at 1553 GMT.

7/24/06

PAD

CIMEL clock is 6 seconds slow.  Reset CIMEL clock at 1523 GMT.  RTS at 1538 GMT.

7/17/06

PAD

CIMEL clock is 1 min. 6 sec. slow.  Reset CIMEL clock at 1516 GMT.  RTS at 1531 GMT.

7/12/06

PAD

CIMEL clock is 4 seconds slow.  Reset clock at 1834 GMT.  RTS at 1849 GMT.

7/10/06

PAD

CIMEL clock is 5 seconds slow.  Reset CIMEL clock at 1826 GMT.  RTS at 1841 GMT.

7/5/06

PAD

CIMEL clock is 5 seconds slow.  Reset clock at 1449 GMT.  RTS at 1504 GMT.

6/30/06

PAD

CIMEL clock is 5 seconds slow.  Reset clock at 1523 GMT.  RTS at 1538 GMT.

6/23/06

PAD

CIMEL clock is 5 seconds slow.  Reset clock at 1455 GMT.  RTS at 1510 GMT.

6/19/06

PAD

CIMEL clock is 8 seconds slow.  Reset clock at 1628 GMT.  RTS at 1643 GMT.

6/12/06

PAD

Instrument status message from mentor states no CIMEL battery voltage.  CIMEL internal battery voltage is 5.7V.  External battery is 14.0V.  Found CIMEL had been set with auto mode off.  Reset CIMEL clock and changed to auto mode.

6/8/06

PAD

CIMEL clock is 11 minutes and 56 seconds slow.  Reset clock at 1508 GMT.  External battery and CIMEL battery voltage checks out.  Troubleshooting at this time.  RTS at 1523 GMT.

6/5/06

PAD

CIMEL clock is 6 seconds slow.  Reset clock at 1552 GMT.  RTS at 1607 GMT.

5/30/06

PAD

Daily report for instrument states no data received for instrument.  At 1428 GMT, found CIMEL was not set in auto mode.  Set clock on CIMEL and changed to auto mode.  RTS at 1443 GMT.

5/26/06

PAD

CIMEL clock is 4 seconds slow.  GOES clock is 4 seconds slow.  Reset CIMEL clock at 1557 GMT.  Reset GOES clock at 1558 GMT.  RTS at 1613 GMT.

5/22/06

PAD

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1829 GMT.  RTS at 1844 GMT.

5/18/06

PAD

CIMEL clock is 5 seconds slow.  Reset CIMEL clock at 2009 GMT.  RTS at 2024 GMT.

5/12/06

PAD

CIMEL clock is 3 seconds slow.  Reset clock at 1518 GMT.  RTS at 1533 GMT.

5/9/06

PAD

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1504 GMT.  RTS at 1519 GMT.

5/5/06

PAD

CIMEL clock is 3 seconds slow.  Reset CIMEL clock at 1518 GMT.  RTS at 1533 GMT.

5/1/06

PAD

CIMEL clock is 5 seconds slow.  Reset CIMEL clock at 1758 GMT.  RTS at 1813 GMT.

4/25/06

PAD

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1541 GMT.  RTS at 1556 GMT.

4/21/06

PAD

CIMEL clock is 3 seconds slow.  Reset CIMEL clock at 1508 GMT.  RTS at 1523 GMT.

4/18/06

PAD

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1511 GMT.  RTS at 1526 GMT.

4/14/06

PAD

CSPHOT clock is 5 seconds slow.  Reset CIMEL clock at 1507 GMT.  RTS at 1522 GMT.

4/10/06

PAD

CIMEL clock is 5 seconds slow.  Reset CIMEL clock at 1528 GMT.  RTS at 1543 GMT.

3/31/06

KT

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1626 GMT.  RTS at 1641 GMT.

3/27/06

PAD

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1649 GMT.  RTS at 1704 GMT.

3/22/06

PAD

CIMEL clock is 4 seconds slow.  Reset clock at 1600 GMT.  RTS at 1615 GMT.

3/7/06

PAD

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1608 GMT.  RTS at 1623 GMT.

2/16/06

PAD

CIMEL clock is 4 seconds slow.  Adjusted clock at 1748 GMT.  RTS at 1803 GMT.

2/10/06

PAD

CIMEL clock is 4 seconds slow.  Adjusted CIMEL clock at 2041 GMT.  RTS at 2056 GMT.

2/6/06

PAD

CIMEL clock is 5 seconds slow.  Reset CIMEL clock at 1601 GMT.  RTS at 1616 GMT.       

1/30/06

PAD

CIMEL is 5 seconds slow.  Reset CIMEL clock at 1725 GMT.  RTS at 1736 GMT.

1/23/06

PAD

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1556 GMT.  RTS at 1611 GMT.

1/17/06

PAD

CIMEL clock is 6 seconds slow.  Reset CIMEL clock at 1558 GMT.  RTS at 1613 GMT.

1/9/06

PAD

CIMEL clock is 5 seconds slow.  Reset CIMEL clock at 1624 GMT.  RTS at 1639 GMT.

12/27/05

PAD

CIMEL clock is 6 seconds slow.  Reset CIMEL clock at 1702 GMT.  RTS at 1717 GMT.

12/19/05

PAD

CIMEL clock is 5 seconds slow.  Reset CIMEL clock at 1749 GMT.  RTS at 1804 GMT.

12/12/05

PAD

CIMEL clock is 6 seconds slow.  Reset CIMEL clock at 1646 GMT.

12/5/05

PAD

CIMEL clock is 5 seconds slow.  Reset CIMEL clock at 1634 GMT.  RTS at 1649 GMT.

11/28/05

PAD

CIMEL clock is 5 seconds slow.  Reset CIMEL clock at 1610 GMT.  RTS at 1625 GMT.

11/21/05

PAD

CIMEL clock is 5 seconds slow.   Adjusted CIMEL clock at 1600 GMT.  RTS at 1615 GMT.

11/16/05

MR

The CSPHOT VITEL battery is dropping below normal.  R/R the VITEL battery at 1928 GMT.  RTS at 1943 GMT.

11/15/05

PAD

CIMEL clock is 6 seconds slow.   Adjusted CIMEL clock at 1626 GMT.  RTS at 1641 GMT.

11/8/05

PAD

CIMEL clock is 5 seconds slow.  Adjusted CIMEL clock at 1615 GMT.  Observed day was set as 28th.  Changed date to 8th.  RTS at 1630 GMT.

11/2/05

MR

The *CSPHOT controller and sensor need to be changed out for calibration.  Removed instrument # 252 and installed instrument control box and sensor #99.  Powered up the instrument and set the clock and checked the settings.  RTS at 1950 GMT.   *Old – 252  *New - 99

10/31/05

PAD

CIMEL clock is 2 minutes 53 seconds slow.  Reset CIMEL clock at 1917 GMT.

10/4/05

PAD

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1634 GMT.  RTS at 1649 GMT.

9/15/05

PAD

GOES clock is 26 sec. slow.  CIMEL clock is 2 min. 29 sec. slow.  Reset GOES clock at 1548 GMT and CIMEL clock at 1549 GMT.  RTS at 1601 GMT.

9/15/05

CAM

CSPHOT transmissions have been successful since replacing the battery on 9/14.

9/14/05

CAM

The CSPHOT transmissions are still failing following the antenna change.  Checked all transmitter settings as per vendor and observed as correct.  Noticed during the transmission that the battery voltage reported 9.3V, even though the battery has been measuring around 13.6V everyday, and even does now when not transmitting.  Vendor suspects the battery capacity may be weak and not allowing transmission.  Disconnected Vitel power cable at 2000 GMT and replaced the battery.  Restored power and set clock at 2015 GMT.  Monitored the battery with a DVM during the next transmission at 2030 GMT.  Battery measured 14.6V before transmission started.  When the Vitel displayed showed transmission on, the DVM reading started fluctuating between 13V and 14.5V.  The Vitel display went blank for about 2 minutes, with the DVM reading fluctuating the whole time.  After about two minutes, the DVM reading stayed steady at 14.6V and the Vitel display showed the next transmission time of 2130.  Will check CSPHOT status page first thing in the morning to see if transmission was successful.

9/14/05

CAM

CSPHOT transmissions are still failing after power cycling the transmitter on 9/13.  Instrument vendor suggested replacing the Vitel antenna.  Powered the Vitel off at 1645 GMT and replaced *antenna.  Restored power and set Vitel clock at 1715 GMT.  During the 1730 GMT transmission, the Vitel showed transmission on for about 1 second and then displayed finished.  Notified vendor of change and they will let us know if the transmissions are working.   *Old – 2T97071362  *New - 171334

9/13/05

CAM

The CSPHOT has not been transmitting files since 5/6/05.  Power cycled the Vitel transmitter at 1821 GMT.  Set date and time at 1822 GMT.  Will check CSPHOT status page when it updates in a few hours to see if transmissions are successful. 

9/12/05

PAD

CIMEL clock is 1 min. 20 sec. slow.  Adjusted CIMEL clock at 1457 GMT.  RTS at 1512 GMT.

9/7/05

PAD

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1410 GMT.  RTS at 1421 GMT.

8/15/05

PAD

CIMEL clock is 15 min. 10 sec. slow.  Battery voltage is 13.6V.  Reset CIMEL clock at 1603 GMT.  RTS at 1618 GMT.

7/28/05

PAD

CIMEL clock is 4 seconds slow.  Reset CIMEL clock at 1515 GMT.  RTS at 1530 GMT.

7/18/05

PAD

Instrument mentor requests Vitel transmitter installed.  Installed Vitel transmitter and set clock.  RTS at 1844 GMT.

7/5/05

PAD

CIMEL clock is 4 seconds slow.  Adjusted CIMEL clock at 1557 GMT.  RTS at 1612 GMT.

6/15/05

PAD

CIMEL clock is 4 seconds slow.  Adjusted CIMEL clock at 1521 GMT.  RTS at 1536 GMT.

5/26/05

MR

The instrument mentor has notified me that the transmitter has stopped communicating with the satellite.  T/S showed that although the transmitters clock was running, the instrument thought the time was not set.  Reset the time and date and activated the transmitter.  RTS at 1545 GMT.

5/9/05

PAD

CIMEL clock is 3 seconds slow.  Reset CIMEL clock at 1455 GMT.  RTS at 1510 GMT.

5/6/05

MR

The new battery charger has arrived and needs to be installed.  Installed the new battery charger in the system.  Power the system up at 2045 GMT.  RTS at 2100 GMT.

5/2/05

PAD

CSPHOT battery is showing 19.9 V.  Disconnected power supply and battery reads 12.2 V.  Removed power supply and disconnected power and communication cable to CIMEL.  Contacted instrument mentor for replacement power supply.  RTS at 1755 GMT.

12/28/04

PAD

At 2053 GMT, GOES clock shows 2035:27 GMT.  Reset GOES clock at 2054 GMT.  RTS at 2109 GMT.

12/23/04

PAD

CIMEL clock is 2 sec. slow.  Reset CIMEL clock at 1646 GMT.  RTS at 1701 GMT.

12/20/04

PAD

CIMEL clock is 4 sec. slow.  Reset CIMEL clock at 1552 GMT.  RTS at 1607 GMT. 

12/15/04

PAD

CIMEL clock is 4 sec. slow.  Reset CIMEL clock at 1546 GMT.  RTS at 1601 GMT.

12/10/04

PAD

CIMEL clock is 3 sec. slow.  Reset CIMEL at 1605 GMT.  RTS at 1620 GMT.

12/6/04

PAD

CIMEL clock is 5 sec. slow.  Reset CIMEL clock at 1538 GMT.  RTS at 1553 GMT.

11/29/04

PAD

CIMEL clock is 6 sec. slow.  Reset CIMEL clock at 1636 GMT.  RTS at 1651 GMT.

11/22/04

PAD

CIMEL clock is 4 sec. slow.  Adjusted CIMEL clock at 1551 GMT.  RTS at 1606 GMT.

11/17/04

PAD

CIMEL clock is 4 sec. slow.  Adjusted CIMEL clock at 1604 GMT.  RTS at 1619 GMT.

11/12/04

PAD

CIMEL clock is 4 sec. slow.  Reset CIMEL clock at 1937 GMT.  RTS at 1952 GMT.

11/8/04

PAD

CIMEL clock is 4 sec. slow.  Reset CIMEL clock at 1612 GMT.  RTS at 1622 GMT.

11/4/04

PAD

Instrument will not align with sun.  Vertical park position is approximately 10 deg. off vertical.  Realigned instrument.  Instrument will not return to same position in park consistently.  Instrument will not align with sun consistently.   Instrument mentor has been contacted.  Troubleshooting at this time. RTS at 1708 GMT.

11/4/04

PAD

CIMEL clock is 5 sec. slow.  Adjusted CIMEL clock at 1623 GMT.  RTS at 1638 GMT.

10/29/04

PAD

CIMEL clock is 3 sec. slow.  Reset CIMEL clock at 1458 GMT.  RTS at 1513 GMT.

10/26/04

PAD

CIMEL clock is 5 sec. slow.  Reset CIMEL clock at 1606 GMT.  RTS at 1621 GMT.

10/22/04

PAD

CIMEL clock is 4 sec. slow.  Reset CIMEL clock at 1513 GMT.  RTS at 1528 GMT.

10/18/04

PAD

CIMEL clock is 5 sec. slow.  Adjusted CIMEL clock at 1446 GMT.  RTS at 1501 GMT.

10/12/04

PAD

CIMEL clock is 4 sec. slow.  Reset CIMEL clock at 1927 GMT.  RTS at 1942 GMT.

10/8/04

PAD

CIMEL clock is 4 sec. slow.  Reset CIMEL clock at 1434 GMT.  RTS at 1449 GMT.

10/4/04

PAD

CIMEL clock is 5 sec. slow.  Reset CIMEL clock at 1442 GMT.  RTS at 1457 GMT.

9/29/04

PAD

GOES clock is 5 min. 6 sec. slow.  CIMEL clock is 4 min. 51 sec. slow.  Reset GOES clock at 1506:30 GMT.  Reset CIMEL clock at 1513 GMT.

9/28/04

MR

The instrument mentor has sent a replacement CIMEL controller and sensor head and it needs to be changed out with the old units.  Shut down the system at 1931 GMT.  R/R the *CIMEL controller and **sensor head.  Reset the time and date.  RTS at 2000 GMT.   *Old -0194992 *New - 0305924    **Old - 0104993   **New - 9707-174 

9/28/04

PAD

CIMEL clock is 3 sec. slow.  Reset CIMEL clock at 1427 GMT.  RTS at 1442 GMT.

8/30/04

PAD

CIMEL clock is 3 sec. slow.  Adjusted CIMEL clock at 1552 GMT.  RTS at 1607 GMT.

8/13/04

MR

The instrument alignment seems to be off.  Performed an alignment procedure as per instruction manual.  Adjusted the base by 1 inch to the west.  RTS at 1900 GMT.

8/12/04

MR

The instrument mentor wants the data cable checked and reseated.  Checked and reseated the data cable and notified the mentor.  He said he would ship us another data cable due to the CIMEL controller does not seem to be communicating with the transmitter.  RTS at 1515 GMT.

8/9/04

PAD

Solar alignment is off.  Troubleshooting to verify date is correct.  Reset CIMEL clock at 1450 GMT.  RTS at 1505 GMT.

8/4/04

MR

The instrument mentor reports that there is no data being received from the CSPHOT CIMEL controller and wants the connections and the controller checked.  The controller seemed to be working ok.  The data cable between the transmitter and the controller checks good.  Reseated the connections and notified the mentor.  RTS at 1800 GMT.

7/15/04

PAD

CIMEL clock is 3 sec. slow.  Reset CIMEL clock at 1522 GMT.  RTS at 1537 GMT.

6/25/04

PAD

CIMEL clock is 4 sec. slow.  Adjusted CIMEL clock at 1528 GMT.  RTS at 1543 GMT.

6/2/04

PAD

CIMEL clock is 4 sec. slow.  Reset CIMEL clock at 1527 GMT.  RTS at 1542 GMT.

4/2/04

MR

The base of the instrument needs to be leveled slightly.  Releveled the base and realigned the instrument.  Status OK.  RTS at 1640 GMT.

2/20/04

MR

The CSPHOT is not holding alignment.  T/S found that the sensor head was slightly set too far back in the bracket which caused the weight to not be distributed right during a scan.  Shifted the sensor slightly forward to balance the weight.  Sensor was then able to scan properly.  RTS at 2039.

2/19/04

MR

The instrument mentor wants the CIMEL controller and sensor exchanged with the one he has sent.  Changed out the *CIMEL controller, sensor and cable as per mentors request.  RTS at 1715 GMT.  *Old – 98   *New - 252 

2/13/04

PAD

During PM rounds, CIMEL clock was observed 4 sec. slow.  Reset CIMEL clock at 1614 GMT.  RTS at 1629 GMT.

2/9/04

PAD

During PM rounds, CIMEL clock was observed 7 sec. slow.  Reset CIMEL clock at 1606 GMT.  RTS at 1621 GMT.

2/3/04

PAD

CIMEL clock is 5 sec. slow.  Reset CIMEL clock at 2006 GMT.  RTS at 2021 GMT.

1/30/04

PAD

CSPHOT CIMEL clock is 5 sec. slow.  Reset CIMEL clock at 1719 GMT.  RTS at 1734 GMT.

1/26/04

PAD

During PM rounds, CIMEL clock was observed 5 sec. slow.  Reset CIMEL clock at 1628 GMT.  RTS at 1643 GMT.

1/20/04

PAD

During PM rounds, CIMEL clock was observed 6 sec. slow.  Reset CIMEL clock at 1725 GMT.  RTS at 1750 GMT.

1/14/04

PAD

During PM rounds, CIMEL clock was observed 5 sec. slow.  Reset CIMEL clock at 1600 GMT.  RTS at 1615 GMT.

1/9/04

MR

Morning rounds found that the motor mount on the CSPHOT has become loose and needs to be tightened.  Tightened the motor mount.  Checked the alignment and it checked good.  RTS at 1645 GMT.

1/9/04

PAD

During PM rounds, CIMEL clock was observed 5 sec. slow.  Reset CIMEL clock at 1545 GMT.  RTS at 1600 GMT.

1/8/04

PAD

CSPHOT has not aligned with sun since 1/5/04.  Found that country in parameter settings had inadvertently been changed to 4.  Reset country in parameters to 2 at 1634 GMT.  CSPHOT will align with sun.  Notified instrument mentor.  RTS at 1649 GMT.

01/05/04

PAD

During PM rounds, CIMEL clock was observed 6 sec. slow.  Reset CIMEL clock at 1730 GMT.  RTS at 1745 GMT.

12/29/03

PAD

During PM rounds, CIMEL clock was observed 6 sec. slow.  Reset CIMEL clock at 1901 GMT.  RTS at 1916 GMT.

12/22/03

PAD

During PM rounds, CIMEL clock was observed 6 sec. slow.  Reset CIMEL clock at 1728 GMT.  RTS at 1743 GMT.

12/15/03

PAD

During PM rounds, CIMEL clock was observed as 6 sec. slow.  Reset CIMEL clock at 1649 GMT.  RTS at 1659 GMT.

12/10/03

PAD

During PM rounds on 12/10/03 at 1725, CSPHOT was found out of alignment.  At 1926 GMT, Mike Rainwater and I checked the alignment.  Unplugged the interface to the CIMEL and rechecked alignment.  Checked all values and time on CIMEL.  Found time was 26 sec. off.  Reset CIMEL clock and checked alignment.  Notified instrument mentor that alignment is off.  Troubleshooting at this time.  RTS at 1946 GMT.

12/8/03

PAD

CSPHOT CIMEL time was observed 6 sec. slow during PM rounds.  Reset CIMEL clock at 1929 GMT.  RTS at 1939 GMT.

12/1/03

PAD

During PM rounds, CIMEL clock was observed 5 sec. slow.  Reset CIMEL clock at 1634 GMT.  RTS at 1649 GMT.

11/24/03

PAD

CIMEL clock was observed 5 sec. slow during PM rounds.  Reset CIMEL clock at 2102 GMT.  RTS at 2117 GMT.

11/17/03

PAD

During PM rounds, CIMEL clock was observed as 5 sec. slow.  Adjusted CIMEL clock at 1734 GMT.  RTS at 1749 GMT.

11/10/03

PAD

CSPHOT CIMEL clock was observed 5 sec. slow during PM rounds.  Reset CIMEL clock at 1616 GMT.  RTS at 1631 GMT.

11/4/03

PAD

CIMEL clock time was observed 5 seconds slow during PM rounds.  Reset CIMEL clock at 1727 GMT.  RTS at 1742 GMT.

10/28/03

PAD

At 1440 GMT, CIMEL clock was observed 5 seconds slow.  Reset CIMEL clock at 1545 GMT.  RTS at 1600 GMT.

10/21/03

PAD

During PM rounds, CIMEL clock was observed 5 seconds slow.  Reset CIMEL clock at 1449 GMT.  RTS at 1504 GMT.

10/14/03

PAD

At 1452 GMT, the GOES clock was 9 sec. slow.  At 1455 GMT, the CIMEL clock was 10 sec. slow.  At 1502 GMT, adjusted the GOES and CIMEL clocks to GMT time.  RTS at 1517 GMT.

10/9/03

KC

CSPHOT was not in proper alignment.  Aligned CSPHOT at 1505 GMT.  RTS at 1522 GMT.

9/22/03

KC

CSPHOT was going to the park position in a sluggish manner. Unplugged and replaced the Photometer plug at 1514 GMT. RTS at 1522 GMT.

9/16/03

KC

Cimel clock was 5 seconds behind GMT clock. Reset Cimel clock at 15:00:00 GMT. RTS at 1515 GMT.

9/3/03

KC

Mentor said the date was not accurate on the Cimel. Reset date, time, and realigned the Cimel at 1448 GMT. RTS at1503 GMT.

8/26/03

MR

The instrument mentor has requested that the physical alignment be checked on the CSPHOT.  Checked the vertical alignment and it was slightly off.  Adjusted the vertical alignment according to his procedures.  Checked the horizontal alignment and it was OK.  RTS at 1500 GMT.

6/30/03

MR

The instrument mentor wants the satellite information changed and the antenna direction turned so that the instrument can use a different satellite.  Changed the VITEL transmitter parameters and direction of the antenna.

6/19/03

JEM

At 1342 GMT, the CIMEL time was observed 9 seconds slow.  Reset CIMEL clock to correct time @1343 GMT.  RTS @1355 GMT.

6/13/03

PAD

Instrument level bubble observed slightly off center.  Leveled instrument then checked alignment.

5/14/03

MR

The instrument mentor wants the internal battery removed from the battery charger and placed on a solar panel.  Installed the solar panel the mentor sent.  Unplugged the internal battery from the battery charger and plugged it into the solar panel.  RTS at 1915 GMT.

4/22/03

ZW

The mentor requested that the date be checked on the CSPHOT. To check the date was not part of the daily, or weekly check. Checked the date as per the mentor request. Found that the year was 2002. Changed the year to 2003. Status OK. RTS 1430 GMT. 

4/9/03

JEM

The CSPHOT is not aligning properly with the sun.  Received the proper latitude and longitude settings from them mentor and they need to be verified.  The following settings were checked and found to be incorrect.  The settings were also changed to the proper values.

Lat mn      +2452    changed to +2196

Lon. HH    W4        changed to W6

Lon. MM   51         changed to 29

Lon. SS      33         changed to 38

After these settings were changed and saved, and alignment check was performed.  Alignment was found to be correct at this point.  RTS @1955 GMT.

3/28/03

MR

The instrument mentor has sent the replacement CSPHOT and needs to be installed.  Installed the new CSPHOT *sensor and **controller.  RTS at 1845 GMT.  *New - 9905-940  **New - 9905-938

3/19/03

JEM

The collimator tubes were observed pointing straight up.  Tried to place the CSPHOT into the parked position @1544 GMT.  No response was received from the CSHPOT.  Manually moved the collimator tubes to point downward so the water could drain out.  Measured battery voltages and the small 12V battery was reading 0.3 VDC.  The large 12V battery was reading 13.0VDC.  Removed the face of the control box to check the internal battery.  As the face was being removed, the smell of burnt electronic components was observed.  After removing the face, a significant number of components were found to be burnt and melted.  Removed the control box and also the sensor at the mentor’s request so they could be shipped to him.  System is down until replacements are received.  RTS @1700

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

12/24/02

JEM

Ice has built up on the collimator tubes and the motor, limiting the movement of the CSPHOT.  Removed all ice from the collimator tubes and the motor so the CSPHOT could move freely @1620 GMT.  RTS @1630 GMT.

12/4/02

JEM

A significant amount of ice has accumulated on the CSPHOT and may be limiting its movement.  Removed as much ice as possible from the CSPHOT.  The CSPHOT moves freely without hanging.  RTS @1655 GMT.

12/2/02

MR

The instrument mentor is requesting that the battery and the battery charger be changed out with the one he has sent.  Changed out the large battery and the charger.  RTS at 1600 GMT.

10/10/02

MR

 The charging system seems to have burned some of the wires leading to and from the battery. T/S found that the old charger that was left in the system shorted out and burned the power wires.  Removed the old battery charger and reconnected and reinstalled the new battery charger.  R/R the power wires to the system.  System came up and started taking data.  Reset the time and date on the VITEL transmitter.  RTS at 1600 GMT.

10/08/02

MR

The instrument mentor wants the battery charger replaced in the instrument due to it not charging the battery fully.  He wants to monitor the system for a couple of days before making this a permanent change. Installed a new battery charger in the system.  Will leave the old charger in the system until it is determined that this change will be permanent.  RTS at 2015 GMT.

9/21/01

MR

The mentor is on site performing an instrument change out and upgrade.  He removed the old *CSPHOT and installed a new unit.  *Old - 98  *New - 252

9/18/02

JEM

The mentor wants to remove the Photovoltaic Charge Control to get the battery voltage to raise from 11.5VDC to above 12VDC.  Disconnected power @1405 GMT.  Removed the Photovoltaic Charge Control and reconnected wires.  Restored power @1410 GMT.  The battery now reads 12.02 VDC.  Status OK.  RTS @1420 GMT.

9/12/02

JEM

Returned to the CSPHOT to reinstall the internal and external batteries @1945 GMT.  Reinstalled the batteries @1950 GMT.  Verified that all the settings were still correct.  Set the Cimel time @ 19:54:00.  Set the Vitel time @19:57:50.  CSPHOT status returned to normal.  RTS @2000 GMT.

9/12/02

JEM

The CSPHOT is not functioning.  The CSPHOTs internal battery voltage has dropped to near 3VDC.  This is to low and will not allow the CSPHOT to run.  Removed the internal battery and put it on a charger.  Will reinstall the battery once it is fully charged.  Did not RTS.

9/11/02

MR

The instrument mentor wants us to check the external batteries and put them on charge due to the time drifting.  Removed the external batteries from the unit and placed them on charge overnight.  Will install the batteries tomorrow morning.

8/5/02

MR

It is time to change out the *CIMEL sensor and controller for calibrations.  The instrument mentor has sent the new units and needs them installed.  Removed the old instruments and installed the new instruments.  Checked the settings of the new instrument and set the time and date.  Tested the controller by sending the sensor to track the sun.  The instrument checked ok.  RTS at 1545 GMT.  *Old - 252   *New - 99

6/14/02

MR

Power restored.  System operational.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT.  The instrument is shut down at this time.

5/9/02

MR

The instrument mentors needs the connections checked and the time and date checked.  Reset the time on the CIMEL clock at 1903 GMT.  Checked the time on the VITEL clock an it checked OK.  Checked the connections and they seemed tight.

10/29/01

MR

The CSPHOT charge regulator has arrived from the mentor and needs to be installed.  Installed the charge regulator.  Further T/S found that one of the solar panels has gone bad.  Due to no solar panels available but AC power is, I replaced the solar panels with a 2.5 amp AC adapter.

10/22/01

MR

The transmitter (GOES) battery is dead.  T/S found that the voltage into the charging regulator was ok but the output voltage was 0 volts.  notified the mentor and requested a new charging regulator.  No further CM at this time.

9/28/01

MR

Installed the CSPHOT #2 transmitter, antenna, battery, solar panels and connectors.  Left the transmitter unplugged due to not being able to reach the mentor.  Will call on Monday 10/01/01 to see if there are any settings we need to change before plugging it in.

7/9/01

MR

Morning rounds found that the large 12 volt battery was dead and only reading 5.8 volts. T/S found that the battery charger had tripped the internal circuit breaker.  Reset the breaker.  After about 1 min the breaker tripped again.  Suspected the battery has gone bad.  Removed the battery and tried to charge it using a larger battery charger.  After 1 hour on the charger the battery would not hold a charge and the voltage was still 5.8 volts.  R/R the battery.  Placed the new battery on a charger to top off the battery.  installed the battery at 1800 GMT.  Reset the system time.  RTS at 1815 GMT.

5/01/01

MR

The mentor wants the voltage readings before and during transmission times. The instrument has not been sending any readable files. Checked the voltage of the battery at 1525. It was 13.30 volts. Checked the voltage reading at 1527 (1 min into transmission) and the voltage reading was 13.28. Checked the voltage reading after transmission and it had climbed up to 13.29 volts. Recorded the readings and sent them to the mentor. Power cycled the transmitter due to it may have been hung up transmitting garbage. Reset the CIMEL and VITEL time. Will await further instructions.

3/20/01

MR

After consulting with Scott Smith from Brookhaven National Lab, he wants the antenna position changed. The Vitel is set up to transmit to GOES west but the antenna is pointing to GOES east. This is thought to be the problem of so many missed messages. Reset the antenna position to an elevation of 30 and an azimuth of 233 at 1720 GMT.

11/01/00

MR

The mentor called this morning and requested that the transmitter be power cycled due to no transmissions being received. Power cycled the transmitter at 1530 GMT. RTS at 1545 GMT.

10/23/00

JEM

The Mentor indicated that Friday the CSPHOT clock was of by 7 minutes. While performing the PM on the CSPHOT the time was observed to be off. Although the GOES time was only off by 34 seconds and the CIMEL time was only off by 33 seconds. When the time was checked the actual time was 15:34:00, the GOES time was 15:33:26, and the CIMEL time was 15:33:27. Set both times and when the times were checked again the actual time was 15:41:00, the GOES time was 15:41:01, and the CIMEL time was 15:41:00. RTS @1542 GMT.

10/3/00

MR

The instrument mentor notified us that the EPROMS needed to be exchanged between the old CSPHOT controller and the new controller. Removed the old EPROMS from controller # 99. Reinstalled the EPROMS in controller #98. Reset the time and date. Checked that all the settings where O.K. Tested the unit and it tested O.K. Communications with the DCP was O.K. RTS at 1530 GMT.

9/27/00

MR

The new *CSPHOT has come in and needs to be exchanged with the old unit. Changed the unit out and found out that the new one does not have the program in it to run and communicate with the DCP. Notified the mentor and awaiting callback. *Old - 99 *New - 98

8/30/00

MR

Reset the system time at 1359 GMT. RTS at 1405.

5/5/00

MR

The CSPHOT was hung up and in a horizontal position. Using the software I placed the unit in the park position and notified the mentor via e-mail. RTS at 1445.

5/2/00

MR

The mentor called and needs the transmitter power cycled. Power cycled the transmitter at 1500 GMT. RTS 1515.

4/17/00

MR

The mentor wants me to call Mr Nadair. Called Mr. Nadair and he stepped me through several tests and checked that all the parameters were set correctly. During one of the checks he said it looked as if a chip inside the sensor head needed to be replaced and will send us a new one with instructions on how to replace it. No further CM performed at this time. Checked and reset the time on both the Cimel and VITEL and placed the unit back in auto mode. No further CM at this time

4/17/00

MR

Reset the time and power cycled the transmitter at 1400 due to no transmissions received.

4/14/00

MR

The replacement transmitter came in today. Installed the transmitter at 520 GMT. Reset the clocks and reset the settings on the VITEL transmitter.

3/30/00

KT

Mentor emailed that there was still no data transmitting. At 1913 GMT power cycled the DCP once again and set the time. RTS at 1930 GMT.

3/29/00

KT

Mentor request that the instrument be power cycled. At 1904 GMT power cycled instrument. At 1914 GMT reset both CIMEL and VITEL times. RTS at 1915 GMT.

3/28/00

KT

Mentor emailed to have the DCP time reset. At 1907 GMT reset both times. RTS at 1923 GMT.

3/27/00

MR

The CSPHOT is not sending files. The unit seemed to be operating O.K. Power cycled the transmitter and reset the time. Status O.K.

3/21/00

MR

The mentor indicates that the CSPHOT has not taken any triplet readings since 3/15/00. The photometer seemed like it was out of alignment. Tried to send the unit to the park position and it would not move. Power cycled the photometer and the transmitter. Again tried to send the unit to the park position and this time it went. Checked the level and it checked O.K. Reset the CIMEL and VITEL time on the transmitter. RTS at 2130.

3/6/00

MR

Power cycled the transmitter due to it not sending data. Reset the time and date. RTS at 1830.

3/3/00

MR

The system clocks are off. Reset both clocks at 1500. RTS at 1515.

2/25/2000

ZW

The mentor requested the satellite transmitter be rebooted. Power cycled the satellite transmitter at 1800 GMT. Status OK. RTS 1815 GMT.

2/24/00

KT

Both the Cimel and Vitel times needed to be adjusted. At 1920 GMT adjusted both the Cimel and Vitel clocks. RTS at 1929 GMT.

2/1/00

MR

Power cycled the transmitter and reset the system clocks. RTS at 1414.

1/28/00

MR

Power cycled the Transmitter. Reset the system clocks. RTS at 1543.

1/25/00

MR

The DAPS download report was showing the CIMEL and VITEL clocks to be off. Reset both clocks. RTS at 1615.

1/19/00

MR

Power cycled the transmitter at 1815 GMT. Reset the system time. Status O.K.

11/29/99

MR

Power cycled the CIMEL controller at 1530 GMT. Reset the time and RTS at 1545.

11/5/99

CAM

CSPHOT instrument mentor indicated that the DCP clock was slow by 34 seconds. Checked the CIMEL and DCP times and found they were both slow by nearly 34 seconds. Set the CIMEL time at 1829 GMT. Set the DCP time at 1834 GMT. RTS at 1835 GMT.

11/2/99

MR

The time was off and 20 files where missing. Power cycled the CIMEL controller and reset the time. RTS at 2150.

10/21/99

KT

The CIMEL and VITEL times needed to be reset. At 1730 GMT in sync with CIMEL time. At 1742 GMT in sync with VITEL time. RTS at 1742 GMT. CIMEL time- 17:42:12 VITEL time- 17:42:09 REAL time- 17:42:12

10/19/99

KT

At 1740 GMT the CSPHOT needed to be re-leveled. At 1742 GMT the instrument was re-leveled. RTS at 1751 GMT.

9/29/99

KT

Instrument mentor reported loss of transmitting data from 9/23/99 to present. At 1950 GMT checked the battery voltages for CIMEL and VITEL which were within operating parameters. At 2000 GMT disconnected the power from the VITEL transmitter. At 2002 GMT the power was reconnected. The CIMEL and VITEL times needed to be adjusted. At 2016 GMT final time correction.        CIMEL time- 20:16:42 VITEL time- 20:16:40. Status ok. RTS at 2022 GMT.

9/23/99

GRS

The Mentor indicated that the Cimel and Goes times are both off. A check of the instrument verified the time problem. Reset both the Cimel and Goes times from 1535 to 1601 GMT. Both times are within two seconds of UCT from Denver Colorado.

9/22/99

KT

At 1656 GMT during rounds the CSPHOT was facing due west and was not moving. At 1927 GMT while observing the instrument the CSPHOT was going through the sample mode and at 1937 GMT after the sample mode was done the CSPHOT returned to the home position. Status ok at this time. RTS at 1940 GMT.

9/21/99

GRS

The instrument mentor indicated that he has not been receiving files from the instrument. Performed some T/S checks and found no problems. Power cycled the Cimel transmitter at 1545 GMT. Will wait and see if this fixes the problem.

8/2/99

MR

Replaced the CIMEL controller batteries with a battery out of a PS-12 Power supply. Asked Mike Ward to reorder the CIMEL batteries. RTS at 1650.

7/28/99

MR

Repositioned the cables due to them binding. RTS at 1530.

7/8/99

MR

Power cycled the transmitter.

7/2/99

MR

Power cycled the transmitter and the controller at 2100. RTS 215.

6/3/99

MW

The CSPHOT stopped sending files starting with the 1546 GMT transmission on 6/2/99. Power cycled the DCP transmitter and reset the GOES and Cimel times at 1900 GMT. The instrument resumed data transmission starting with the 1946 GMT transmission on 6/3/99.

4/30/99

MR

Changed out the EPROMS. Changed the lat and long values to match the old values in the old controller. RTS at 1922.

4/29/99

MR

Changed out the *photometer and **controller. RTS at 1900. *Old - 9710-256 *New - 9707-174 **Old - 9710-255 **New - 9707-173

4/1/99

MW

Today's DAPS download report indicated that once again, no files were being received from the instrument. The Vitel transmitter battery measured 2.2 Vdc and the charger power switch was in the off position again. Flipped the power switch on and measured the battery voltage at 10+ Vdc. Had to reset the DCP time at 1541 GMT. Not sure if transmission will be successful this time with such a low battery voltage. Will generate a PR for a new battery today. Troubleshooting on the charger problem continues.

3/30/99

MW

The CSPHOT DAPS download data report sent to site ops on 3/30/99 indicated that no files had been received from the instrument since 3/29/99 at 0600 GMT. The Vitel transmitter battery measured only 3.3 Vdc. The battery charger power switch was in the off position. The charger is currently set to charge 24 hours a day. Flipped the chargers power switch to the on position at 1935 GMT. The battery came up to 11+ Vdc. The next DAPS download report indicated that files were received after performing the aforementioned corrective maintenance.

3/3/99

MW

The CSPHOT had not transmitted a data file since 0600 GMT on 3/2/99. Took the instrument out of auto mode at 1630 GMT. Changed opmode to off and power cycled the Vitel transmitter. Reset the DCP time and changed opmode back to self. Placed the instrument back in auto mode at 1636 GMT. Will not know if this corrected the problem until the next data dump supplied by the instrument mentor.

2/12/99

ZW

The Mentor for the CSPHOT called to report that the CSPHOT was not transmitting. Site Ops power cycled the Vitel Satellite Transmitter at 1650 GMT. Site Ops also set the DCP (GOES) time, and the controller time. Status ok. RTS 1730 GMT.

2/9/99  

ZW

Mr. Rangshai requested Site Ops check to see if the CSPHOT cable was binding at certain times during sky scans. The CSPHOT cable would catch on the corner of the base plate. Site Ops used a grinder to smooth off the northeast, and northwest corners of the square base plate. Status ok. RTS 2100 GMT.

1/19/99

MW

The CSPHOT download report indicated no files received since 1/14/99. Took the instrument out of auto mode at 1950GMT. Changed 'OpMode' to 'off'. Power cycled the Vitel transmitter. Changed 'OpMode' to 'self' and placed the instrument in auto mode. RTS at 2000 GMT.

12/14/98

MK

The CSPHOT was not sending any files via GOES transmission. Took the instrument out of auto mode and disabled GOES transmission. Power cycled the DCP at 1604 GMT. Reset the DCP time after the power cycle. Enabled GOES transmission and placed the photometer back in auto mode. Returned to service at 1625 GMT.

11/23/98

MW

The instrument mentor indicated that no files were being received from the CSPHOT and he requested that the DCP be power cycled. Took the instrument out of auto mode and disabled GOES transmission. Power cycled the DCP at 1605 GMT. Reset the DCP time after the power cycle. Enabled GOES transmission and placed the photometer back in auto mode. Returned to service at 1630 GMT.

11/18/98

MW

The photometer will not line up correctly with the sun when given the 'gosun' command. The instrument mentor is aware of this and we are currently trying to isolate the problem. The photometer does not seem to be able to find the correct home position and when the 'gosun' command is given, the photometer alignment is off by a corresponding amount. Contacted the instrument mentor and requested that he provide site operations with the correct list of parameters for the CSPHOT so that they can be compared to the parameters which are currently in effect.

11/10/98

SS

Daily checks revealed that the sun alignment is off. Checked the bubble alignment the bubble is dead center. Attempted to align it manually but found that the instrument would not align to the sun correctly the azimuth is good but the elevation will not come up to the needed angle. Checked the configuration and found that it looks correct. Will continue to trouble shoot on 11/12/98 with Mike Ward and the instrument mentor.

11/4/98

MW     

The instrument mentor requested that the current photometer be removed and sent in for calibration. He sent a replacement photometer and asked that it be installed. Powered the system down at 1750 GMT. R/R *photometer and **Cimel control box. Checked all parameters and reset date and time of the Cimel and DCP. Set OpMode to 'Self' and placed the system back in 'Auto' mode at 1845 GMT. RTS. *Old - 9707-174 *New - 9710-256 **Old - 9797-173 **New - 9710-255

10/26/98

MK

The DCP clock was off . At 1632 attempted to reset clock but was unsuccessful due to memory error. Reset the DCP clock at 1850 GMT. RTS at 1900 GMT.

10/6/98

MW

Checked the CSPHOT batteries today and they have recharged fully. Placed the photometer in the park position at 1647 GMT. Turned GOES transmission back on and placed the unit in auto mode at 1648 GMT. RTS.

10/5/98

MW

The CSPHOT photometer was not in the proper park position. The batteries had been drained over the weekend after a power outage and didn't have enough power to move the robot arm to the proper park position. Took the instrument out of auto mode and turned off GOES transmission at 1635 GMT. Will let the batteries charge up overnight.

9/30/98

MW

The CSPHOT was not sending any files via GOES transmission and previous reports from the mentor indicated that the Cimel and DCP clocks were out of tolerance. Checked the time of the Cimel and DCP. Both clocks were within 2 seconds of actual WWV time. Took the instrument out of auto mode and disabled GOES transmission at 1513 GMT. Power cycled the Vitel transmitter. Reset the DCP time, re-enabled Goes transmission, and placed the unit back in auto mode. RTS at 1525 GMT.

9/11/98

SS

WO #3040 mentor requests that the DCP be power cycled due to no files received since 9/9/98. Cycled power to the DCP. Time needed set due to power cycle. Set time and checked settings. No problems found. RTS at 1630 GMT.

9/8/98

SS

The instrument is out of level and solar alignment. leveled instrument and checked solar alignment. Solar alignment was out of tolerance. Aligned to solar and had to re-level again. This process was repeated several times before satisfactory results were achieved. RTS at 1600 GMT.

8/27/98

SS

Daily checks revealed that the level and solar alignment was slightly off. Leveled the base and checked solar alignment is now good. RTS at 1530 GMT.

8/24/98

SS

The bubble and solar alignment were off. Leveled the instrument and aligned to the sun at 1543 GMT.

8/12/98

CAM

Unable to precisely check alignment due to cloud coverage, however the alignment visually appears to be off. Will verify alignment when conditions allow.

7/31/98

MW

The CSPHOT stopped sending files. Power cycled the GOES transmitter at 1715 GMT. File collection restored. RTS at 1730 GMT.

7/17/98

CAM

Instrument mentor indicated that no files had been received from CSPHOT since 1646 GMT on 7/14/98. Measured the DCP battery voltage to be 14.2 VDC. Power cycled the Vitel transmitter at 1932 GMT. Reset the DCP time at 1933 GMT. Will continue to monitor transmission status to see if this solves the problem.

7/10/98

CAM

Instrument mentor advised that correcting the charger problem on 7/9 fixed the CSPHOT instrument, but there now appears to be a problem with the Vitel transmitter. Measured the Vitel battery voltage to be 13.8V. Power cycled the transmitter at 1400 GMT and set the clock. RTS at 1410 GMT. Mentor later advised that power cycling the transmitter solved the problem.

7/9/98

MR

The CIMEL batteries had discharged. Examined the battery chargers and they seemed to be working. Observed that the timers for the chargers were 4 hour behind the current time and that the on and off switches were not in the right position which was allowing the chargers to charge at night instead of during the day. Reset the timers and the CSPHOT started to properly work at 1745. RTS at 1805.

6/30/98

MK

The CSPHOT is not transmitting properly. Rebooted the GTM at 1613 GMT, reset the GTM time and RTS at 1635.

6/25/98

CAM

Instrument mentor indicated that the CSPHOT has not been transmitting since 1347 GMT on 6/22/98. Mentor requested the Vitel transmitter be power cycled. Removed power from Vitel transmitter at 1808 GMT. Restored power at 1810 GMT. Took system out of auto mode and turned the DCP transmission off at 1828 GMT. Reset Vitel time at 1829 GMT. RTS at 1830 GMT. Will check with mentor to see if CSPHOT resumes data transmissions this afternoon.

6/19/98

SS

Instrument is intermittently stopping. R/R charger timer to the Cimel. RTS at 1515 GMT.

6/17/98

SS

Mike Ward requested that the time be verified with WWV. Set wrist watch with WWV and verified time on the instrument. Results WWV 15:16:00 Vitel 15:16:03 Cimel 15:16:01. No adjustments made.

6/12/98

CAM

Continued T/S of maintenance earlier today: Checked the Cimel battery voltage on 6/12 and found that it was only 10.0 VDC. Checked the output of the AC/DC charger and found it was not charging the battery. Further T/S revealed the Radio Shack timer that had been placed in the charging circuit had failed. The timer was not turning the outer dial which contains the 'On' and 'Off' tabs. As a result the AC/DC charger was not being activated to charge the battery. Placed the timer in a continous on mode for now until a new timer can be purchased. The Cimel battery voltage increased to 10.5 VDC and is now charging. The CSPHOT returned to normal operating mode at 1930 GMT on 6/12/98. RTS status O.K.

6/12/98

CAM

During daily PM checks on 6/9/98 the CSPHOT was observed to be out of position between samples. The robot was pointing towards the sun constantly rather than returning to the home position. Took the instrument out of auto mode and issued the park command at . Robot returned to park position and the instrument was then returned to auto mode. The robot was later observed to be staying in the sun position again. Instrument mentor was out of town. Left message for mentor to contact Site Ops. Mentor called on 6/12 and advised to power cycled the system. Power cycled system per mentors instructions at 1450 GMT on 6/12. The robot again remained in the sun position after the first sample. Phoned mentor and left message detailing what had occurred. T/S still in progress.

6/10/98

KT

Re-leveled the bubble level at 1705 GMT. RTS at 1713 GMT.

6/4/98

CAM

During daily PM checks the CSPHOT bubble level was observed to be out of the center circle. No adjustment made at this time. Adjusting the level may also affect the alignment. Sky conditions are currently overcast which would not allow proper alignment check following level adjustment. Will level when sky conditions are more favorable.

5/11/98

MW

The Vitel satellite transmitter parameters would not stay in the proper configuration. The parameters would change randomly every time an attempt was made to set the parameters to the proper configuration. Also, a 'DCP error #11' and 'DCP error #13' were seen when viewing the parameters. R/R the *Vitel transmitter 1715-1730 GMT. *Old - VX9707196 *New - VX9707197.

5/5/98

MW

The DCP time was reset to 00:00:00. Took the instrument out of 'auto' mode at 1626 GMT. Attempted to take the DCP out of transmission mode but the CIMEL displayed the message 'Saving Error!' when switching 'OpMode' from 'self' to 'off'. The DCP will not allow the adjustment of the time setting without stopping the GOES transmission first so the time could not be adjusted. Removed power from the DCP to reset it at 1629 GMT. Repeated the attempt to take the DCP out of transmission mode once again but was unsuccessful. Checked under INFD for errors and found 6 errors indicated. The DCP was displaying 'DCP Error #11' and 'DCP Error #13' alternately. Cannot change the DCP time so the instrument was left out of 'auto' mode until further troubleshooting can be done.

4/14/98

KT

The sun tracker was not aligned with the sun. At 1445 GMT aligned the tracker up with the sun. RTS at 1510 GMT.

4/7/98

MW

1510-1530 GMT - Took the system out of 'auto' mode and disabled GOES transmission. Removed the Vitel from the system enclosure (left all connectors in place). Measured voltage between the failsafe terminal connector on the Vitel and the negative terminal of the 12V battery at 0Vdc. Inserted a copper wire into the failsafe terminal connector and shorted it to the positive terminal of the 12V battery for 10 seconds to reset the failsafe as per mentor request. Re-installed the Vitel into the system enclosure. Re-started GOES transmission and placed the instrument back in 'auto' mode. Checked Cimel and Vitel system times. Times OK. Checked the following parameters at mentor's request: "DCPTRANSM=YES", "DCP 1H=YES", "DCP LOADT=15". Changed "DCP LOADT=15" to "DCP LOADT=30" at 1932 GMT. Verified that data was being transmitted from the Cimel to the Vitel - at 2030 GMT: "Nb Bytes=1267", at 2050 GMT: "Nb Bytes=0". Checked all connections. T/S still in progress.

4/6/98

MW

Installed an AC charger on the Cimel batteries and installed timers in both the Cimel and Vitel battery charging circuits 2100-2130 GMT. The batteries are set to charge between 1200-2200 GMT. The robot is now functioning normally. RTS at 2130 GMT.

4/2/98

MW

The CSPHOT instrument mentor indicated that no data was being collected from the instrument via GOES satellite transmission. Checked GOES transmission status and the OpMode was set to "self". Took the system out of auto mode and disabled GOES transmission and power cycled the Vitel. Reset the Vitel time and re-enabled GOES transmission and placed the system back in auto mode. Checked the times of the Cimel and Vitel and saw the message "DCP error #13. Measured 14.8V across the Vitel 12V battery. All Vitel parameters look OK. Unplugged charger from Vitel battery (2110 GMT) at mentor's request to ensure that the battery wasn't overcharging. T/S still in progress.

4/2/98

MW

The CSPHOT robot arm was not functioning properly. It would not lock on to the sun and it acted as if the motors didn't have enough power to position the photometer in the park or sun positions. Checked output of solar panel at 15Vdc. Checked the voltage level of the two 6V Cimel batteries in series and measured 5Vdc. Suspect the robot drained the batteries over the weekend of 3/28-3/29 due to overcast conditions which prevented the solar charger from charging the batteries. The instrument probably hasn't worked since then.

4/1/98

MW

Noticed the photometer was not pointing at the sun. It was only elevated about 22 degrees from the park position, pointing out in the field South of the instrument. Took the instrument out of "Auto" mode at 1650 GMT. Instructed the instrument to go to the "Park" position. As the photometer was going towards the park position, the motors sounded louder than normal as if they were straining. After the photometer was in the park position, the instrument was given the "GOSUN" instruction. Once again the motors sounded as if they were straining and the instrument did not make a smooth transition from the "Park" to the "GOSUN" position. The photometer rose only about 22 degrees from the park position as it was attempting to point at the sun. Adjusted the position of the detector head by loosening the holding bracket and sliding it up so the bracket was clamped around the bottom edge when it was in the park position. This enabled the detector to rise a little higher but still not high enough to point at the sun. Tried several mounting positions for the detector head but could not find a "sweet" spot where it would easily go to both the "sun" and the "park" positions correctly. The zenith motor seems too weak to correctly move the detector into the proper position. T/S still in progress.

3/31/98

MW

The CSPHOT instrument mentor requested that the Vitel transmitter be replaced since no data has been transmitted from the instrument since it was installed. Took the instrument out of auto mode at 1615 GMT. Stopped the GOES transmission. R/R *Vitel control box. Reset the transmitter time at 1624 GMT. Verified all Vitel parameters and restarted GOES transmission. Placed the unit back in auto mode at 1630 GMT. RTS. *Old - VX9707197 *New - VX9707196

3/17/98

MW/SS

The CSPHOT has not been sending files since it was installed on March 13th. The instrument mentor requested that the GOES transmitter be reset and the time updated. Took the instrument out of AUTO mode at 1927 GMT. Stopped the GOES transmission and power cycled the Vitel box. Reset the time on the Vitel box at 1930 GMT. Restarted GOES transmission and placed the instrument back in AUTO mode at 1931 GMT

3/11/98

Mentor

Instrument installed by mentor (Rangasayi Halthore). Site Ops received training.